OpenStudy (anonymous):

Help here please: Find the exact value of the following, show all work: a. a. sin^2(75°

OpenStudy (experimentx):

you know cos2A = ??

OpenStudy (anonymous):

cos (30) = sqrt(3)/2 that?

OpenStudy (experimentx):

5:30 am here ... actually i was planning to sleep.

OpenStudy (experimentx):

\[\cos(2A) = (cosA)^2 - (sinA)^2\] \[\cos(2A) = 1 -2 (sinA)^2\] \[1- cos(2A) = 2(sinA)^2\] \[(1- cos(2A))/2 = (sinA)^2\]

OpenStudy (anonymous):

Oh. Im so sorry! I didnt know! ok then, I will ask for help elsewhere, Thanks for your help today! Good night:)!

OpenStudy (experimentx):

I will stay for a while ... sin^2(75) = (1-cos150)/2

OpenStudy (anonymous):

Oh ok:) so I can find the exact value by applying the formula you provided from above?

OpenStudy (anonymous):

is the problem just to find sin^2(75) ?

OpenStudy (anonymous):

Its actually, sin^2(75°) / sin 45° cos^2 30°

OpenStudy (anonymous):

 like that?